### v1

# Gradient-free neural topology optimization

Abstract
Gradient-free optimizers allow for tackling problems regardless of the smoothness or differentiability of their objective function, but they require many more iterations to converge when compared to gradient-based algorithms. This has made them unviable for topology optimization due to the high computational cost per iteration and high dimensionality of these problems. We propose a pre-trained neural reparameterization strategy that leads to at least one order of magnitude decrease in iteration count when optimizing the designs in latent space, as opposed to the conventional approach without latent reparameterization. We demonstrate this via extensive computational experiments in- and out-of-distribution with the training data. Although gradient-based topology optimization is still more efficient for differentiable problems, such as compliance optimization of structures, we believe this work will open up a new path for problems where gradient information is not readily available (e.g. fracture).
Keywords Gradient-free optimization  $\cdot$ Generative deep learning  $\cdot$ Latent optimization  $\cdot$ Topology Optimization
1 Introduction
Topology optimization plays an important role in the design of structures. Historically, method development has been focused on relatively simple problems, such as compliance minimization that are linear and differentiable [1, 2, 3]. While these problems are of high importance, many practical scenarios are governed by non-differentiable, non-linear objectives that are much more challenging to optimize for [4, 5, 6]. In principle, gradient-free optimization algorithms could be a promising approach for solving such problems because they can be applied to any objective. However, they suffer from a major limitation: they require several orders of magnitude more objective evaluations (simulations) than gradient-based optimizers to converge to a solution.
Gradient-free optimizers update the solution by sampling and comparing the performance of trial solutions. At each iteration, the objective needs to be evaluated for the whole population of samples, using expensive simulations (usually FEM). Moreover, gradient-free optimizers suffer from the curse of dimensionality, i.e. the cost of optimization grows exponentially with the number of design variables [7, 8]. This is especially nefarious to topology optimization, as the number of design variables is typically very large rendering these optimizers unfeasible [1, 8].
Considering these limitations, it comes as no surprise that gradient-free optimization is not well motivated for most topology optimization problems [8]. However, there are scenarios where the use of gradient-based methods might be challenging, namely in cases where the gradients are difficult to compute or where the objective function is noisy or discontinuous [4, 9]. In principle, gradient-free methods would be a logic solution in these contexts, but their slow convergence detracts from their use [8].
Contributions
In this article we address the large rift in performance between gradient-free and gradient-based topology optimization for problems where the latter is orders of magnitude more efficient than the former. Therefore, we consider thousands of topology design problems where gradient-based optimizers severely outperform the best gradient-free algorithms and show that an appropriate machine learning strategy combining training (offline) and neural reparameterization (online) closes the performance gap by an order of magnitude.
Compared to the state-of-the-art latent optimization of topology with gradient-free methods [10], our work explicitly targets the issue of scalability of gradient-free optimization. We show that the appropriate choice of the model architecture leads to significant gains in performance, compared to the conventional approach (without latent space reparameterization), but also compared to more standard architectures.
We test our approach on out-of-distribution examples, i.e. examples that were not used in training and that include characteristics that were not explored in training. Furthermore, we deploy our model on an optimization problem with different physics and significantly different design features, to demonstrate promising generalization performance. Notwithstanding, we do not claim that the proposed approach can outperform gradient-based methods in convex or nearly convex objective functions, as occurs in the compliance optimization problems that we consider in this work. However, we believe that this research is a first step towards solving more challenging problems in the future, involving plasticity, fracture, and fatigue.
2 Related work
2.1 Limitations of gradient-based approaches
In many optimization problems the objective function is differentiable, such as optimization of compliance, and the derivatives can be obtained analytically or with automatic differentiation [11, 12, 1, 13]. In case the function is non-differentiable, so far, the most widespread approach is to linearize the problem such that it can still be solved with a gradient-based optimizer, in particular using the adjoint method [14, 6, 5, 15, 16]. Although relatively successful, this approach comes with certain limitations, as highlighted for example in [4], where the authors use the adjoint formulation of the topology optimization problem to postpone fracture. The presented derivation of the adjoint assumes that the objective function is smooth, however, this is not the case for fracture, where even the smallest changes in the design can have a critical impact on the performance (specifically, these changes can completely change the crack trajectory). As highlighted by the authors, the effects cannot be easily remedied by fixing the step size of the descent algorithm or by regularization. In such a case, when the objective function is not well behaved, discontinuous, with multiple local minima, and challenging to regularize, gradient-free methods could be a promising alternative [8, 4].
Several gradient-free algorithms were explored in topology optimization, including Genetic Algorithms [17, 18], Artificial Immune Algorithms [19], Ant Colonies [20, 21], Particle Swarms [22], Simulated Annealing [23], Harmony Search [24], Differential Evolution [25] among others [8]. These approaches, however, received strong criticism in the community [8] primarily due to their inferior efficiency, as opposed to gradient-based methods. Gradient-free algorithms suffer from the curse of dimensionality [8, 7], which means that the cost of optimization increases approximately exponentially with the number of problem dimensions. In standard topology optimization approaches [11], where the design variables correspond to element-wise densities, the dimensionality of even small ’toy-examples’ is in practice too large for gradient-free algorithms, leading to prohibitive costs (in order of 10,000 - 100,000 objective function evaluations, as opposed to the order of 10 – 100 evaluations for gradient-based methods [8]).
2.2 Latent-space optimization
The curse of dimensionality of gradient-free optimizers can be addressed by reparameterizing the optimization problem into a lower-dimensional latent space [26, 27, 28]. This dimensionality reduction can be well accommodated using generative deep learning models, such as generative adversarial networks (GAN) [29], variational autoencoders [30], and diffusion models [31, 32]. In the context of latent space optimization, variational autoencoders are a particularly common choice [33, 34].
Optimizing in latent space of a VAE was demonstrated in context of topology optimization [35, 36, 10]. Sato et al [35] and Gladstone et al. [36] explored latent optimization with gradient-based approach using a surrogate model – a neural network trained to predict the property of interest from the latent representation. The method takes advantage of the differentiability of the surrogate - as the property of interest can be minimized by backpropagating through the surrogate. In both cases, the results demonstrate that VAE is capable of representing new designs, that outperform those seen in the training. The disadvantage of this approach, however, is the need for data, required for training the surrogate.
Guo et al. [10] explore topology optimization in the latent space of a VAE with different optimization algorithms, including a gradient-free genetic algorithm. Unlike other works, the training of the VAE was augmented with an additional style-transfer procedure, to improve the quality of the generated designs. The training dataset comprised designs optimized for thermal compliance using the SIMP method [11], but the model was then used to solve a different problem - a multiobjective optimization, minimizing the maximum temperature and maximizing the power density. Similar to other works [35, 36], the method was able to synthesize novel designs, different to those seen in the training dataset. Most importantly, the results demonstrate that the pre-trained model can be successfully deployed on a different optimization problem, given the designs have similar features. Nevertheless, in the presented experiments the optimization process with the gradient-free method (Genetic Algorithm) required an order of 30,000 FEM evaluations, indicating relatively limited improvement as compared to the gradient-free optimization without reparameterization into the latent space [8].
2.3 Machine learning in topology optimization
In the context of topology optimization, generative machine learning models have been primarily used to map boundary conditions to final designs [37, 2, 38]. Different generative models were explored in such a setup, including GANs [39, 40] and diffusion models [41]. While demonstrating competitive performance, these approaches are often criticized for their limited generalization [37, 2], as they remain applicable only to the problem where they were trained on. In order to apply such a model to a new type of design problem (e.g. different objective, different physics), the model needs to be retrained on a dataset corresponding to the new problem. Generating the dataset, however, requires the use of conventional optimization methods and simulations.
Another relevant stream of work is neural reparameterization [12, 42, 13, 43, 2], where the neural network is used to implicitly represent the designed structure. Instead of optimizing the element-wise density values, the optimizer adjusts the trainable parameters of a neural network which outputs the density field. This approach proved beneficial in gradient-based approaches, where the weights can be adjusted by back-propagating the gradients of the density field (design sensitivities) [12]. Neural reparameterization was not explored in a gradient-free optimization scenario. One reason is the relatively high number of parameters of the neural networks, ranging from 100s [43] to 100,000s [44], making most network architectures unfeasible to optimize with gradient-free algorithms.
3 Methods
We propose a latent space optimization approach following two steps, as shown on the schematic in Figure 1. First, we train a generative model to reparameterize the topology designs into a lower-dimensional latent representation. Then, in the second step, the latent space is explored by the gradient-free optimizer, which attempts to find a design minimizing the objective. The optimizer controls the latent vector which parameterizes the design. This vector is decoded into a physical design using the pre-trained generative model, and the design is evaluated using the objective function (e.g. compliance). During one iteration of optimization the optimizer generates a population of trial latent vectors, which are subsequently evaluated, and based on their associated objective values, the optimizer updates the population according to its internal update rules, converging to a better solution.
(a) Schematic of a variational autoencoder (VAE). In the first step of our method we train the generative model - a VAE to reparameterize topology designs into latent space. Here $\mu$ and $\sigma$ are mean and variance vectors in latent space, $\epsilon$ is the noise vector sampled from a multivariate Gaussian, and $z$ is a latent variable.
(b) Schematic of latent space optimization process with gradient-free optimizer. In the second step of our method, we use a trained generative model, to optimize the designs using latent space representation of the generative model. Note that the latent vector $z$ is no longer stochastic, and it is fully controlled and updated ($z_{new}$) by the optimizer, based on the objective values obtained from the simulator.
Figure 1: Schematics of a) the generative model and b) the latent space optimization setup used in our proposed methods.
The key assumption of our approach is that the generative model learns compact representations of features occurring in structural designs, but these designs, in principle, are not necessarily bound to any particular objective function. We observe that the designs optimized for different objectives (e.g. compliance, minimum mass, or compliant mechanism design) oftentimes share the same set of visual features (e.g. compare the designs optimized for fracture to those optimized for standard compliance in [4]). This observation can be verified using quantitative feature similarity metrics such as Learned Perceptual Image Patch similarity [45] (for details, see the appendix). Based on that, we assume that the generator trained to capture visual features can be deployed to optimize different objectives without the need for retraining, and is not strictly limited to the problems used for generating the training dataset. This is a crucial advantage over state-of-the-art approaches exploring generative models for topology optimization.
3.1 Generative Model
We propose the use of a new VAE variant called Latent Bernoulli Autoencoder (LBAE) because it uses the Bernoulli distribution for modeling the latent space [46]. Although this architecture was proposed in a different context, it has been shown to achieve superior performance compared to conventional VAE architectures in terms of quality of the generated samples (FID score). We conjectured that using a Bernoulli distribution for the latent space (zeros and ones), instead of a continuous distribution such as a Gaussian, is particularly adequate for the case of topology optimization where the aim is to decide whether each discrete element should contain material or not. We carefully compare the LBAE to a conventional VAE architecture and searched for the best hyperparameters for a large class of problems, in an attempt to provide consider each method at their best performance [10]. Furthermore, we include gradient-based results and different gradient-free optimizers to establish clear performance metrics and clarify the usefulness of the proposed method.
Conventionally, VAE provides a generative model that captures the distribution of data $\mathcal{D}=\left\{\mathbf{x}^{(1)},\mathbf{x}^{(2)},...,\mathbf{x}^{(N)}\right\}$, by assuming a joint distribution $p(\mathbf{x},\mathbf{z})$ using some hidden (latent) variable $\mathbf{z}$, such that given $\mathbf{z}$, it can generate samples of $\mathbf{x}$. In practice, this is achieved by using two neural networks: the encoder, which approximates the posterior $p(\mathbf{z}|\mathbf{x})\approx q_{\phi}(\mathbf{z}|\mathbf{x})$, (since in general $p(\mathbf{z}|\mathbf{x})$ is intractable), and the decoder, which captures $p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})$. The two networks are trained simultaneously by maximizing the evidence lower bound, as shown in Equation 1.
$L_{\mathbf{\theta},\mathbf{\phi}}=\operatorname{\mathbb{E}}_{q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\frac{p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})p(\mathbf{z})}{q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})}\right]=\operatorname{\mathbb{E}}_{q(\mathbf{z}|\mathbf{x})}\left[\log p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})\right]-D_{KL}\left(q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})||p(\mathbf{z})\right)$ (1)
The first term corresponds to the reconstruction loss, and the second term corresponds to the Kullback-Leibler divergence, which regularizes the latent variable and aims to drive the parameterized distribution towards the assumed prior. In a conventional VAE, the prior $p(\mathbf{z})$ is assumed to be a multivariate Gaussian, while in LBAE the latent space is assumed to be a multivariate Bernoulli distribution. Unlike in VAE, however, in LBAE the Bernoulli distribution of the latent variable is imposed via network architecture and its specific sampling routine, rather than by an explicit regularization term (for details see [46]). Therefore, the second term from Equation 1 is not present in the LBAE loss.
The training dataset $x\sim X$ in our case corresponds to images representing different topology designs. Once trained, the autoencoder is then used for generating new designs by sampling $z$ from the latent space. In this way, the designs can be represented using a lower dimensional latent vector.
Architecture
The LBAE architecture that we use is based on the model proposed by Fajtl et al. [46], adapted to our dataset by adjusting the number of layers and convolutional filters. According to the authors [46], modeling latent space with Bernoulli distribution produces sharper images, compared to traditional VAE, and allows for smoother interpolation in the latent space. We hypothesize that these properties are crucial for optimization, allowing for easier traversing of the loss landscape by the optimizer. Furthermore, unlike in the baseline VAE, LBAE uses residual layers. Combined with the Bernoulli distribution of the latent variable, we observed that this architecture indeed is capable of producing outputs with finer detail and reduced blurriness, addressing one of the main shortcomings of VAE [2]. The baseline VAE architecture used in our case is a modified model of Larsen et al. [47]. The inputs and outputs are single-channel, grayscale representations of the designs of resolution 64x64.
For both LBAE and VAE, one of the most important parameters is the dimensionality of the latent space, which balances the training and the optimization performance. While increasing the latent space dimensionality makes the training easier, as the compression of information is not as severe, the optimization becomes less efficient due to the curse of dimensionality [8, 7]. To find a balanced latent space dimensionality, we tested several configurations (for details, see the appendix). In the case of VAE, we did not observe a significant difference between the dimensionality 32 and 64, instead, we noticed that the performance was limited by insufficient resolution of finer details in the output (caused by the blurriness intrinsic to the VAE). These limiting factors, however, were to a large extent alleviated in LBAE, where the trade-off due to the latent space dimensionality was much more clear (for details, see the appendix). We found that for LBAE the dimensionality 256 offered the best balance.
Dataset
We use two different datasets of topology designs at train and test time, respectively. The training dataset was generated based on a modified approach of [48] and [41], resulting in 48,872 designs. The boundary conditions were generated as randomly drawn point loads and point supports, using Poisson’s distribution specified in the same way as in [48]. Additionally, in our dataset, we randomly applied a square mask to restrict part of the design domain from applying material (e.g. for an L-shaped beam the mask would be applied in the corner). The mask was applied with 25% probability, and the position of the mask was randomly assigned along one of the edges of the design domain. Volume fractions were drawn using a uniform distribution with a range of [0.12, 0.5]. This dataset was split into two subsets: for training and validation, using a standard 80:20 split.
The testing dataset was used only for monitoring the testing loss, which aims to verify the generalization capability of the model. This dataset was generated using the problems defined in [12]. The boundary conditions were generated using one of 28 classes of parameterized problems, with the parameters drawn at random from a uniform distribution, which resulted in over 4000 instances of testing problems. Unlike the training problem set, the testing problem classes include distributed loads and supports, therefore they are expected to provide a good insight into the generalization of the model and can be considered out-of-distribution.
The designs in both datasets were obtained using a gradient-based MMA optimizer with the default hyperparameter settings. The final designs were represented as grayscale images of fixed resolution (64x64) with pixel values representing the volume fraction of the material. For each problem, the optimization with MMA [49] was run for 200 iterations. The simulations were set up following the approach of the 88-line code [11], with $E_{0}=1.0$, $E_{min}=1e-9$, $\nu=0.3$ and a standard cone filter with a radius of 2 pixels for regularizing the density field.
3.2 Benchmarking set-up
We develop and benchmark our method considering a standard topology optimization problem, namely compliance minimization with a volume constraint. While this problem can be easily solved using gradient-based approaches, it is not the ultimate application of the proposed method. Nevertheless, it provides a good benchmark for the development of the method because, first, the "ground truth" solutions can be easily obtained using a gradient-based approach, allowing us to verify ’how far’ our method is from a good solution. Secondly, these problems are relatively cheap to evaluate, as opposed to design problems based on more complex physics.
Benchmarking problem sets
To provide statistically meaningful performance measures, we test our optimization framework on two sets of optimization problems. The first set contains 25 in-distribution design problems (sets of boundary conditions). These problems were generated in the same way as the problems seen during the training (i.e. with random point loads and point supports). The second set contains 25 out-of-distribution problems, generated using the approach of Hoyer et al. [12] (the same as the testing dataset). These problems include distributed loads and supports, therefore they provide a better insight into the generalization of the model.
Performance metrics
To quantify the performance of different model and optimizer configurations across the whole problem set, for each problem, we normalize the compliance value with the compliance of the MMA solution (considered as the ’ground truth’), and in this way calculate the relative error. Furthermore, to account for stochastic factors within the optimization process (e.g. initialization of the optimizer), the optimization process for each configuration and each problem is repeated with 15 different random seeds. The performance is evaluated for grayscale and binary (thresholded) designs. For thresholding we apply the simplest strategy from [50] in a post-processing step. In that case, the relative error is calculated with respect to the value obtained for the thresholded MMA design.
Gradient-free optimizer
For gradient-free optimizer, we use Covariance Matrix Adaptation Evolutionary Strategies (CMA-ES) [44, 51] and its Bi-Population variant (BIPOP-CMA-ES) [52]. The CMA-ES and its variants (such as BIPOP-CMA-ES) are currently regarded as state-of-the-art gradient-free optimization algorithms [7]. Additionally, CMA-ES was considered in a similar setting [53]. Notwithstanding, we considered several alternative gradient-free optimizers to ensure that we selected the most favorable one for our problems, namely Particle Swarm Optimization [54], Differential Evolution [55], Evolutionary Strategies meta-optimizer [56], and Simple Genetic Algorithm [57]. However, even with hyperparameter tuning these optimizers exhibited considerably worse performance than CMA-ES within the assigned budget.
Topology optimization set-up
The volume constraint is enforced using a constrained sigmoid transformation, following the approach of Hoyer et al. [12]. For the remaining parts, we use the same set-up as for the MMA baselines - i.e. the cone filter with a radius of 2 pixels, and the material parameters set to $E_{0}=1.0$, $E_{min}=1e-9$ and $\nu=0.3$.
4 Results
We found that while optimizing topology without gradients, parameterizing the problem using the latent space of an autoencoder speeds up the process by at least one order of magnitude, closing the gap to gradient-based optimizers reported in the literature. Due to the choice of neural architecture, we demonstrate that the method shows significant robustness and generality, unlike elsewhere in the literature. These findings are demonstrated in Figure 2, which compares our LBAE and VAE architectures against the conventional baseline approach without reparameterization (pixel parameterization). In this example, the designs were optimized with BIPOP-CMA-ES (for complete results, including other optimizers, see the appendix). The performance is quantified as a fraction of problems (the cumulative probability) vs. the relative error of the final solution with respect to the MMA solution – considered here as the ground truth.
Figure 2: Cumulative probability (fraction of problems) vs relative error w.r.t. the MMA solution, for two variants of latent parameterization models - Variational Autoencoder (VAE, latent space dimensionality: 64), and Latent Bernoulli Autoencoder (LBAE, latent space dimensionality: 256), compared against conventional pixel parameterization. The designs were optimized with a BIPOP-CMA-ES optimizer. Note that the shown distributions do not add up to 1 due to the cut-off at 300%. Best, typical, and worst designs correspond to the best, median, and worst random initializations for each problem.
The results are grouped by problem set (in-distribution, and out-of-distribution problems). To investigate the robustness of our method, we also analyze the results considering only subsets of runs - for each design problem out of 15 randomly initialized optimization runs, we select the best, median, and worst run, to calculate the cumulative probability, as shown in 2 in columns 2-4 respectively. In all cases, the LBAE reparameterization significantly outperforms conventional black-box optimization within the prescribed evaluation budget (2000 FEM simulations), regardless of whether we choose all, best, median, or worst runs. In addition, the worst runs of LBAE remain significantly better than the best runs of the conventional pixel parameterization method, proving the overall robustness of this approach.
As expected, the performance on out-of-distribution problems is worse than on in-distribution problems. At first glance, this can be attributed to the differences between the types of problems seen in the train and test sets. However, worse performance was observed for all the methods, including conventional black-box without reparameterization (done without the use of an autoencoder), which indicates that these out-of-distribution problems are objectively more challenging to optimize. This outcome is particularly encouraging, as it illustrates the difficult task that the proposed method was subjected to.
Comparing LBAE architecture with the VAE, LBAE achieves significantly better performance overall, with the most significant difference visible on the subset of the worst runs, indicating on robustness of our approach. The VAE still achieves a significant margin over the conventional black-box optimization, thus proving the overall benefit of reparameterizing into the latent space, even with a more naive architecture. Nevertheless, the margins between LBAE and VAE indicate that the choice of architecture has a crucial influence on the performance.
Additionally, we find that the performance in terms of relative errors remains relatively intact whether it is evaluated on grayscale or thresholded designs. This is an important finding since historically many generative models struggled with producing representations with a significant presence of gray values, which after thresholding leads to disconnected structures (and thus a significantly deteriorated performance) [2]. However, in our case, the designs obtained with LBAE suffer from the gray-value problem to the same extent as the designs obtained with MMA.
Despite a significant advantage over the conventional pixel parameterization, the final objective values when compared to the gradient-based solutions can still be substantial – some designs show errors on the order of 100%, i.e. the compliance value is double the one found with MMA. For these problems, the imposed budget of 2000 FEM evaluations turns out insufficient to find a good solution. Therefore, unsurprisingly, the designs obtained without gradients remain inferior to those obtained with MMA.
Testing other optimizer configurations (CMA-ES, BIPOP-CMA-ES, with default and customized hyperparameters) led us to the same overall conclusions: latent reparameterization leads to significant improvement over conventional pixel parameterization, and LBAE in particular consistently outperforms the other methods (for details, see the appendix).
Example designs
We investigate further the performance of our method by analyzing example designs.4. Figure 3 shows an example problem L-shaped bracket design problem from the out-of-distribution set. The example is representative of the most common cases in both problem sets – the relative errors achieved with LBAE with respect to the MMA solution are relatively low, with the worst solutions being within 10-30% worse than MMA solutions in terms of the relative error, and the best designs within 5-10%. The designs are visually quite different from the ground truth, and also different from each other - highlighting the stochastic nature of the gradient-free optimization. Note that for LBAE, the thresholding plays in favor of the model, leading to even lower relative error values as compared to the MMA design. In the case of VAE, however, the solutions are significantly oversimplified, and lacking detail, which results in significantly worse performance.
Figure 3: An example of an out-of-distribution problem: classic L-shaped bracket problem, with clamped support along the top edge and a downward point load applied at the right edge.
Figure 4: An example of a challenging out-of-distribution problem: a ’staircase’ design with Z-shaped distributed load, with clam support a the bottom edge.
Figure4 shows an example of a more challenging out-of-distribution problem, with a Z-shaped distributed load, requiring a more intricate arrangement of the material. The variation in performance is relatively large, with the worst LBAE design yielding 132% higher compliance compared to the baseline. The designs generated with VAE achieve orders of magnitude worse performance. As in the previous problem (Figure 3), the designs obtained with VAE are distinctly oversimplified, indicating on limited capacity of the VAE architecture as compared to LBAE.
Interestingly, thresholding affects the design performance more drastically, and inconsistently as compared to the first example in Figure 3. The reason is the connectivity of the designs, and more specifically how it is affected by the thresholding step. Noteworthy, some of the designs appear to perform much better after thresholding e.g. typical design in LBAE outperforms the design of MMA. The reason for that is that the design obtained with MMA does suffer to some extent from intermediate ’gray-value’ densities present at the elements with prescribed external forces. As a result of thresholding, some of these externally loaded elements become void of material, altering the connectivity and increasing significantly the compliance value of the MMA design. In that case, the placement of the material within the elements subject to external loads in the LBAE design turns out more favorable, and robust to thresholding, leading to better performance.
Finally, all designs obtained with the conventional black-box optimization approach with pixel parameterization were so far from convergence, that we did not include them in the plots. This behavior was observed for all the tested problems in both problem sets, which indicates that the budget of 2000 FEM evaluations is simply insufficient for the conventional black-box optimizer to converge. This observation is in good agreement with the values reported in literature [8], stating typical numbers in order of 10,000 - 100,000 FEM iterations until convergence.
Heat Conduction problem
In order to explore the limitations of our model in terms of generalization, we deploy our model (LBAE), without retraining, on a fundamentally different problem – optimization of heat conduction. While optimization for heat conduction can be easily solved with a gradient-based approach, it provides an excellent benchmark for our purpose. Governed by different physics as compared to the structural problems (although it is the same type of Partial Differential Equation), the typical boundary conditions and consequently the solutions are distinctly different. In particular, the design features occurring in thermal problems, are different than those seen in structural compliance optimization (oftentimes slender structures with beam-like elements).
We set up a simple heat-conduction example following that from [58]: there is a uniformly distributed heat source across the square design domain, and a heat sink in the middle of one of the edges, with the length of 0.2 of the edge length. The results obtained with our methods are compared to MMA in Figure 5. Even though our model has never been exposed to thermal problems of any kind, it can produce a solution, with a final objective (1.70) relatively close to that obtained with MMA (1.28). The baseline VAE achieves a better score than LBAE, however the design is less reminiscent of the one obtained with MMA (ground truth).
Figure 5: Thermal compliance problem: comparison of latent optimization with CMAES and Latent Bernoulli Autoencoder (LBAE) vs MMA, unprocessed solution as well as thresholded
Additionally, we consider objectives evaluated on thresholded designs, where the densities are rounded to 0 or 1, to eliminate the gray values, while preserving the volume fraction. Thresholding highlights the challenge of this design problem – the objective is very sensitive to the intermediate density values, making it an interesting benchmark. Due to the uniform heat source, the optimizers attempt to cover as much surface as possible, even at the cost of distributing intermediate densities of material, leading to gray values. As a result, the value of the objective evaluated on the thresholded design is two orders of magnitude larger, as compared to the objective evaluated on the design that has not been thresholded. Comparing the thresholded designs, latent optimization (LBAE with CMAES) leads to much worse performance, with the objective over four times larger as opposed to that obtained with thresholded MMA design. The thresholded design obtained with VAE yields much higher (worse) objective when compared to LBAE.
This problem also offers a good opportunity to demonstrate some advantage of the gradient-free approach – instead of optimizing for the continuous density design (as required by the gradient-based approach), we can easily change the objective to optimize directly for the thresholded design objective. In other words, instead of optimizing for the objective evaluated with continuous density values, and thresholding in a postprocessing step, we can provide the optimizer with objective values evaluated directly on the thresholded designs. Optimizing for this objective leads to a much better loss value, as shown in 6, which drops almost fourfold compared to that in 5, and is now only 20% higher than the thresholded MMA design.
Figure 6: Thermal compliance problem: latent optimization with CMAES and Latent Bernoulli Autoencoder (LBAE), optimized for thresholded objective, compared with MMA optimized for unthresholded objective and, thresholded afterwards, in the postprocessing step.
5 Discussion
Model limitations
We observe that the designs generated by LBAE oftentimes tend to be simpler, than those obtained with MMA (see Figure 3). While in certain scenarios this can be a desired feature (e.g. for manufacturing), it appears that in this case, the simplicity originates from some deficiencies of the generator model, which struggles with producing finer features.
To verify these limitations, we carry out a simple study on the expressivity of the model. We consider several sample designs (obtained with MMA) and pass them through the LBAE, to investigate whether the model has enough expressivity for reconstruction. As shown in Figure 7, the designs reconstructed from the latent encoding provided by the encoder are imperfect and lack the finer features. Next, we take the latent encoding of each design and optimize it (fitting MSE loss to the MMA design, using ADAM optimizer), to verify whether the model has enough expressive capacity. The results of the decoded designs with fitted latent vectors are shown in the bottom row. These designs incorporate finer features, providing more accurate and detailed representations of the MMA target designs. Nevertheless, the blurriness of these finer features is not completely eliminated, indicating expressivity limits of the decoder.
Figure 7: Examples of baseline designs obtained with MMA, and their reconstructions obtained with LBAE. Note that LBAE still exhibits some blurriness wwhen reconstructing finer features.
We speculate that these effects could be partially attributed to the convolutional layers – known to exhibit higher impedance to finer features [59, 60] and the spectral bias of neural networks [60]. In other words, convolutional neural nets struggle more with fitting fine features than coarse ones. This deficiency, however, could be potentially addressed by extending the training, or further adjustments in the architecture. As shown in the presented examples in Figure 3 lack of fine features can have a detrimental influence on the performance of the designs, and thus the overall optimization capabilities of the proposed method. Addressing those limitations might lead to significant improvements in the model’s performance.
Finally, the approach using variational autoencoder architectures based on convolutional neural networks is by design constrained to uniform meshes, and restricted in terms of resolution. Furthermore, we have not tried to solve 3D problems and we do not know how the approach would scale in those scenarios. Further research is required to explore new architectures that tackle these limitations.
Optimization limitations
The second component of the performance of our method comes from the optimization procedure and its limitations. The major point of attention in that regard is the setting of the constraints. Specifically, the volume constraint was imposed following the method of [12], by applying a constrained sigmoid transformation on the output of the convolutional layers of the generator. However, this way of enforcing constraints could potentially hinder latent optimization. During the training, the VAE was exposed to samples with different volume fractions, which resulted not only in different shapes but in different topologies. Therefore, imposing the constraint directly in latent space by restricting to a manifold corresponding to a particular volume fraction could be more beneficial. Finding an alternative way of enforcing the constraint could potentially further improve the performance of the proposed latent optimization method.
Benchmarking
In our benchmarks, while comparing different methods, we used objective values reached at the same number of FEM simulations (simulation budget). This approach gives a fair overview of the total computational cost required for optimization, assuming that FEM comprises the majority of the computational cost in the optimization loop. However, most gradient-free optimizers (including CMA-ES used in this study) rely on evaluating a population (or a batch) of trial designs that can be done in parallel (lowering the total time significantly). This is not the case for gradient-based approaches, where the simulations are strictly sequential. In our set-up, the population sizes are in order 10, which means that the number of iterations (with parallelization – the compute time) is one order of magnitude less than the budget, as shown in the example in Figure 8. Consequently, while considering the parallelization, the gap to gradient-based approaches closes even further.
Figure 8: Comparison of loss curves for an example out of distribution problem. Given that for gradient-free optimization the simulations can be executed in parallel for each population, comparing performance based on iteration is more representative for the actual time (and not the computational cost, as with comparison by FEM evaluations). From this perspective, the gap between gradient-based and gradient-free optimizers closes even further.
6 Conclusion
In this study, we presented a proof of concept framework for optimizing topology designs without gradients in a latent space of a Latent-Bernoulli Autoencoder (LBAE). We found that the proposed method significantly improves the scalability of gradient-free optimization – addressing the major shortcoming of this approach. Our extensive benchmarks indicate that reparameterizing the optimization problem into a lower-dimensional latent space of LBAE leads to dramatic gains in performance, with at least one order of magnitude faster convergence as opposed to the conventional black-box optimization without reparameterization. We demonstrated the robustness of our method, by showing that the performance gains are significant even in the worst-case scenario i.e. considering only the worst runs out of all random initialization. Furthermore, we investigated the generalization performance, by deploying the LBAE model on a completely new problem (thermal compliance optimization) without retraining. Even though the design features desired in this problem are significantly different from those seen in the training dataset, the LBAE-driven optimization converges to a reasonable design, with performance close to that of the gradient-based baseline, proving promising generalization capabilities.
Nevertheless, our implementation still suffers from several limitations. We decided to impose hard constraints (volume fraction), which potentially hinders the optimization performance. We speculate, that imposing constraints directly in the latent space might potentially lead to better optimization. The results indicate a bias towards certain feature length scales in the designs. Generating features over a wider range of length scales, however, is desired for improving the expressivity and variety of the solutions. Finally, our current architecture is limited to uniform grid meshes and a restricted resolution. Convolutional layers may not scale well to 3D applications. Therefore, exploring different architectures such as Graph Neural Networks is expected to lead to even more significant performance gains, minimize the latent space dimension, further improve the scalability, and extend the applicability of latent space optimization. Still, we are surprised by the large performance improvements we found in our work relative to a conventional gradient-free approach to topology optimization, and we find particularly encouraging to show for the first time the generalization ability for out-of-distribution cases.
Acknowledgments
The authors would like to acknowledge that this effort was undertaken in part with the support from the Department of the Navy, Office of Naval Research, award number N00014-23-1-2688. Part of this research was conducted using computational resources and services at the Center for Computation and Visualization, Brown University.
Appendix A Complementary results
Figures 9, 10 and 11 present results for the remaining optimizer configurations: CMA-ES with the default settings, as well as CMA-ES and BIPOP-CMA-ES with customized settings (added variable clipping and initialization, both within the range of [-5, 5]).
Figure 9: Results for different reparameterization models, optimized with CMA-ES optimizer with default settings
Figure 10: Results for different reparameterization models, optimized with CMA-ES optimizer with customized settings
Figure 11: Results for different reparameterization models, optimized with BIPOP-CMA-ES optimizer with customized settings
Appendix B Thresholded designs
Figures 12, 13 and 14 present the results for thresholded designs. The relative error is calculated with respect to the compliance obtained for a thresholded MMA design.
Figure 12: Statistics for the thresholded designs, optimized with CMA-ES with default settings.
Figure 13: Statistics for the thresholded designs, optimized with CMA-ES with custom settings (variable clipping and initialization within the range [-5, 5]).
Figure 14: Statistics for the thresholded designs, optimized with BIPOP-CMA-ES with custom settings (variable clipping and initialization within range [-5, 5]).
B.1 Comparison of optimizers
Figures 15, 16 and 17 compare the performance of 2 different optimizers, with different settings (default and custom). In the custom configuration we apply clipping of the latent variables in the range of [-5, 5], and initialize the latent variable within that range.
Figure 15: Comparison of optimizers using VAE parameterization
Figure 16: Comparison of optimizers using conventional pixel parameterization
Figure 17: Comparison of optimizers using LBAE parameterization
B.2 Optimization with thresholding
The results shown in Figure 18 were obtained using thresholding during optimization, i.e. the objective value was evaluated and optimized on an already thresholded sample design.
Figure 18: Caption
B.3 Correction of the latent vector
To account for the limitations arising from imposing the volume constraint in a hard way, we experiment with a strategy to account for that effect by correcting the latent representation. Given a latent representation, we generate a target design with an imposed volume fraction constraint (sigmoid with constrained mean). Given the same latent representation, we generate a design without the constraint (applying only standard sigmoid activation in the output layer). The latent vector is then optimized by minimizing the MSE loss between the target design (with the volume constraint) and the design without the constraint. The gradients are backpropagated through the decoder network, and the updates to the latent vector are applied using ADAM optimizer, with learning rate 0.1, using 10 iterations. The results are compared with the standard approach, without correction in Figure 19.
Additionally, we experiment with applying tanh activation on the latent variable, before passing it through the decoder network, as a way to emulate the Bernoulli distribution, as experienced during the training. As can be seen in Figure 19, both strategies turn out to deteriorate the performance.
Figure 19: Results for different correction schemes, compared to the default LBAE configuration, optimized using BIPOP-CMA-ES optimizer.
B.4 Convergence rate estimations
Figure 20 presents mean error (averaged over all designs from both datasets, disregarding 10% of the worst runs) vs number of objective evaluations. Although each problem is characterized by a different optimization history, affected e.g. by initialization, the plot gives a qualitative indication of the convergence rates between the 3 different parameterization schemes.
Figure 20: Results for different correction schemes, compared to the default LBAE configuration, optimized using BIPOP-CMA-ES optimizer.
Appendix C Shape similarity argument
Here we present a quantitative argument supporting our assumption that the design similarity could be a reasonable proxy for good performance. In other words, we show that designs optimized for different objectives oftentimes are similar (share the same features). We consider the designs of the MBB-beam reported in work by Desai et al. [4], and compare the design optimized for compliance against designs optimized for fracture with different hyperparameter settings. We quantify the similarity using the Learned Perceptual Image Patch Similarity (LPIPS) metric [45], which provides superior performance for quantifying perceptual similarity by using the feature maps of pre-trained neural networks. The similarity indices for different designs are shown in Figure 21, computed using two different feature map extractors (VGG and Alex-Net).
The compliance design is slightly more dissimilar to the fracture designs - as indicated by a higher LPIPS index. However, we also find a counter-example - the fracture design (extracted from figure 8b in [4]), with a comparably high similarity index. In other words, the fracture design (8b) is as similar to the remaining fracture designs as the compliance design. The findings are consistent regardless of which feature extractor was used for LPIPS calculation.
Figure 21: Learned Perceptual Image Patch Similarity (LPIPS) index, quantifying similarity between the designs reported by Desai et al [4] optimized for compliance and for fracture using different hyperparameters.
Appendix D Latent dimensionality and population size trade-off
Table 1 illustrates the trade-off present while choosing the latent space dimensionality. Models with larger latent dimensionality are easier to train (requiring less compression). On the other hand, larger dimensionality hinders the optimization performance of gradient-free optimizers. In our study, the latent dimensionality of 256 was found to balance these two factors best.
Table 1: Average relative error (wrt. MMA solution) obtained with LBAE models with varying latent space dimensionality, and with different population sizes (CMA-ES).
Population size
Latent
dimension
4 8 16 32
512 1.156 0.575 0.600 0.712
256 1.289 0.507 0.512 0.659
128 3.637 0.817 0.623 0.761
64 15.677 1.151 0.835 1.090
32 2.777 1.709 1.280 1.642
Similarly, the optimal population size (number of samples for an update of the gradient-free optimizer), is also governed by a trade-off. In this case, it is the frequency of the updates (lower population – more frequent updates) against the quality of the updates (higher population – better estimation of the optimal update). In that regard, the population size on the order of 10 samples provides the best performance.
Appendix E Alternative configurations of the baseline VAE model
For the baseline VAE model we tested 4 different configurations: 2 different architecture configurations (see Table 2), and 2 different parameterizations of the latent space. The results are summarized in Figure 22.
Figure 22: Cumulative probability vs relative error (w.r.t. the MMA solution) for different optimization methods, evaluated on smaller testing sets (10 problems each, unlike on the full results), repeated with 10 different random seeds. The distribution does not add up to 1 due to cut-off at 300%
Table 2: Two VAE architecture configurations were used in this study. The decoder uses the same layers but in reversed order, with bilinear upsampling in between the convolutional layers to increase the resolution.
Encoder configuration L Encoder configuration S
convolutional layers (4)
filters: 64, 64, 128, 128
stride 2
kernel size: 4
padding: 1
activation: Leaky ReLU
filters: 32, 32, 64, 64
stride 2
kernel size: 4
padding 1
activation: Leaky ReLU
Dense layer:
input dim: 2048,
output dim: 512,
activation: Leaky ReLU
input dim: 1024,
output dim: 256,
activation: Leaky ReLU
Dense layer:
input dim: 512
output dim: 2x64 ($\mu$ and $\sigma$)
activation: linear
input dim: 256
output dim: 2x32 ($\mu$ and $\sigma$)
activation: linear
Parametrization
For optimization with VAE, we explore two different parameterizations of the latent space: 1) where the latent optimizer controls the whole input to the decoder ($\vec{z}$), and 2) where we utilize the information from the encoder, and the optimizer controls the additive term $\epsilon$, such that the input to the decoder $z=\mu+\sigma\epsilon$, where $\mu$ and $\sigma$ are outputs of the encoder. In this second approach, we feed the best design obtained so far throughout the optimization roll-out, initially starting with a uniformly distributed material. In such a way, the optimizer controls the perturbation of the best design so far, rather than the complete design.
In the results presented in the paper we use the larger architecture (VAE-L) with standard parameterization of the latent space (i.e controlling the full latent vector).